The PH of a solution that is 0.1M HA(1*10^-6 would be

Grade:12th pass

1 Answers

PERUMAL
10 years ago
PH =1/2[PKW-PKA-logc =1/2[14-1-1] =1/2[12] =6 hense, this problem solve.